# On the Characteristics of the Conjugate Function Enabling Effective Dual Decomposition Methods

### v1

@@ I-B Our Contribution
  • Benefits of FGOR on the global consensus problem: we explore how the FGOR characteristics can be exploited when solving the global consensus problem [cf.  (29)] using dual decomposition. In particular, we show that by utilizing FGOR characteristics, we can improve the performance of the standard dual decomposition algorithm not only in terms of the speed of the convergence but also in terms of efficient communication, cf.  § V, Algorithm 3, Lemma 1, and § VI.
  • Lipschitzian properties of the dual function $g$: under less restricted assumptions, we establish that the gradients $\nabla g$ of $g$ are Lipschitz continuous, a more general Lipschitzian property than the existing results, cf.  Appendix B, Proposition 2, Corollary 2.
@@ I-C Notation
Normal font lowercase letters $x$, bold font lowercase letters $\mathbf{x}$, bold font uppercase letters $\mathbf{X}$, and calligraphic font $\mathcal{X}$ represent scalars, vectors, matrices, and sets, respectively. The set of real numbers, set of extended real numbers, set of real $n$-vectors, set of real $m\times n$ matrices, set of positive integers, and set of nonnegative integers are denoted by ${\rm I\!R}$, $\overline{{\rm I\!R}}$, ${\rm I\!R}^{n}$, ${\rm I\!R}^{m\times n}$, $\mathbb{Z}_{+}$, and +$\mathbb{Z}^{0}_{+}$, respectively. The boundary of a set $\mathcal{X}$, the interior of a set $\mathcal{X}$, and the convex hull of a set $\mathcal{X}$ are denoted by $\texttt{bnd}~{}\mathcal{X}$, $\texttt{int}~{}\mathcal{X}$, and $\texttt{con}~{}\mathcal{X}$, respectively. The domain of a function $f:{\rm I\!R}^{n}\to{\rm I\!R}$ is a subset of ${\rm I\!R}^{n}$ and is denoted by $\texttt{dom}~{}f$. The set of all the subgradients of a function $f$ at a point $\mathbf{x}\in{\rm I\!R}^{n}$ is denoted by $\partial f(\mathbf{x})$. The range of a matrix $\mathbf{X}$ is denoted by $R(\mathbf{X})$. For a given matrix $\mathbf{X}$, $\mathbf{X}^{\mbox{\scriptsize T}}$ denotes the matrix transpose. The positive definite cone is denoted by $\mathbb{S}^{n}_{++}$. The $n\times n$ identity matrix and the $n$-vectors with all entries equal to one are denoted by $\mathbf{I}_{n}$ and $\mathbf{1}_{n}$, respectively.
@@ I-D Organization of the Paper
The rest of the paper is organized as follows. In § II, we investigate the FGOR characteristic of the conjugate function. The FGOR characteristic of the dual function is explored in § III. In § IV, the application of FGOR on the dual subgradient algorithm is discussed. Employing FGOR to solve the global consensus problem is discussed in § V. Numerical experiments are presented in § VI. Finally, we conclude the paper in § VII, followed by appendices.
II FGOR Characteristic of the Conjugate $f^{*}$
In this section, we demonstrate an appealing property of the conjugate function $f^{*}$ of $f$ [cf. problem (2)], that is, FGOR, which we will employ to devise our stepsize rules. Roughly speaking, we show that there is a specific region in the domain of $f^{*}$ such that for all points $\nu$ in the region, there is a ray originating from $\nu$ along which the gradients of $f^{*}$ remain constant. Besides, note that the dual function, $g$ is simply a restriction of $-f^{*}$. As a result, further, we show that $g$ directly inherits the FGOR characteristic from $-f^{*}$ if the restriction conforms to certain conditions. Then we leverage the FGOR characteristic of $g$ to design a precursory yet simple and communication-efficient stepsize rule that can be prepended with other existing ones for improved convergences.
@@ II FGOR Characteristic of the Conjugate $f^{*}$
The function $f_{0}$ is lower semicontinuous (lsc), proper, and strictly convex. The set $\mathcal{Y}$ is convex and compact.
Note that from the practical point of view, the compact constraint set is not typically a restriction. For example, it is a common approach in many machine learning problems that impose regularization as a constraint [45, 46]. It can be shown that the problem of convex regularized minimization is equivalent to its corresponding regularization-constrained formulation [47]. Moreover, decision variables pertaining to feasible resources associated to engineering problems are usually finite [41, 48].
@@
By the assumption above, one can exclude cases of infinite gradients at the boundary of $\mathcal{Y}$. The function $l=l_{0}+\delta_{\mathcal{Y}}$ in Example 5 [cf.  Appendix B] is a case where the assumption above breaks since the gradient of the function $l_{0}$ tends to $\infty$ or $\infty$- ∞ as $\bar{y}$ tends to $1$ or $-1$, respectively.
Next, we outline a few remarks that are useful in deriving the FGOR characteristic of $f^{*}$.
Remark 1
@@
$\mathcal{V}=\texttt{con}~{}\{\partial f_{0}(\mathbf{\bar{y}})\ |\ \mathbf{\bar{y}}\in\texttt{bnd}~{}\mathcal{Y}\}$ (3)
Then $\bm{\nu}_{0}\in\texttt{int}~{}\mathcal{V}$ if and only if $\exists~{}\mathbf{y}\in\texttt{int}~{}\mathcal{Y}$ such that $\bm{\nu}_{0}\in\partial f(\mathbf{y})$.
We illustrate the set $\mathcal{V}$ and the idea of Remark 1 using the following example for clarity.
Example 1 (An Illustration of Remark 1)
@@
Then $\partial f_{1}(\mathbf{\bar{y}})$, where $\mathbf{\bar{y}}=[\bar{y}_{1}\ \bar{y}_{2}]^{\mbox{\scriptsize T}}\in\texttt{bnd}~{}\mathcal{Y}_{1}$ is given by
$\partial f_{1}(\mathbf{\bar{y}})=\begin{cases}\displaystyle[4\bar{y}_{1}\ 0]^{\mbox{\scriptsize T}}\ \ \ \,;&\ \text{$\bar{y}_{1}\in[0,1]$, $\bar{y}_{2}=0$}\\ \displaystyle[4\ 4\bar{y}_{2}]^{\mbox{\scriptsize T}}\ \ \ \,;&\ \text{$\bar{y}_{1}=1$, $\bar{y}_{2}\in[-1,0)$}\\ \displaystyle[4\bar{y}_{1}\ -4]^{\mbox{\scriptsize T}};&\ \text{$\bar{y}_{1}\in[0,1)$, $\bar{y}_{2}=-1$}\\ \displaystyle[0\ 4\bar{y}_{2}]^{\mbox{\scriptsize T}}\ \ \ \,;&\ \text{$\bar{y}_{1}=0$, $\bar{y}_{2}\in(-1,0)$}.\end{cases}$
Therefore the set $\mathcal{V}$ associated to $f_{1}$ is given by $\mathcal{V}=\{{\bm{\nu}\in{\rm I\!R}^{2}\ |\ -[0\ 4]^{\mbox{\scriptsize T}}\leq\bm{\nu}\leq[4\ 0]^{\mbox{\scriptsize T}}\}}$, cf.  Figure 1LABEL:sub@subfig:LevelSets-of-f1. Similarly, one can show that the set $\mathcal{V}$ associated to $f_{2}$ is given by the polyhedron with vertices $-[3\ 3]^{\mbox{\scriptsize T}}$, $[-3\ 1]^{\mbox{\scriptsize T}}$, $[-1\ 3]^{\mbox{\scriptsize T}}$, $[3\ 3]^{\mbox{\scriptsize T}}$, $[3\ -1]^{\mbox{\scriptsize T}}$, and $[1\ -3]^{\mbox{\scriptsize T}}$, cf.  Figure 1LABEL:sub@subfig:LevelSets-of-f2. Furthermore, to demonstrate the idea of Remark 1, consider the set $\mathcal{V}$ associated to $f_{1}$ with $\mathbf{u}=[2\,-2]^{\mbox{\scriptsize T}}$ and $\mathbf{v}=[0.5\,-0.5]^{\mbox{\scriptsize T}}$. Clearly, $\mathbf{u}\in\texttt{int}~{}\mathcal{V}$ and $\mathbf{u}=\nabla f_{1}(\mathbf{v})$, where $\mathbf{v}\in\texttt{int}~{}\mathcal{Y}$. Conversely, for $\mathbf{r}=[0.25\,-0.75]^{\mbox{\scriptsize T}}$ and $\mathbf{s}=[1\,-3]^{\mbox{\scriptsize T}}$, $\mathbf{r}\in\texttt{int}~{}\mathcal{Y}$ and $\mathbf{s}=\nabla f_{1}(\mathbf{r})$, where $\mathbf{s}\in\texttt{int}~{}\mathcal{V}$.
@@
Suppose Assumption 1 and Assumption 2 hold. Then for all $\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int~{}}\mathcal{V}$, there exists $\bm{\eta}\in{\rm I\!R}^{n}$ such that for all $\alpha\geq 0$, $\bm{\nu}+\alpha\bm{\eta}\in{\rm I\!R}^{n}\setminus\texttt{int}~{}\mathcal{V}$ and $\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})$ is a constant vector.
@@
Let $\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int}~{}\mathcal{V}$. Then, Remark 1 confirms that $\exists~{}\bar{\mathbf{y}}\in\texttt{bnd}~{}\mathcal{Y}$ such that $\bm{\nu}\in\partial f(\bar{\mathbf{y}})$, i.e.,
@@
where $\bar{\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})$ and $\bm{\eta}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})$, cf.  Remark 3. Now consider the ray
@@
where the equality follows from (4). Since $\bar{\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})$ and $\bm{\eta}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})$, $\bar{\bm{\nu}}+(1+\alpha)\bm{\eta}\in\partial f(\mathbf{\bar{y}})$ for all $\alpha\geq 0$, cf.  Remark 3. Therefore, [42, Proposition 11.3], together with  (5), we have
@@
for all $\alpha\geq 0$. Moreover, since $\mathbf{\bar{y}}\in\texttt{bnd}~{}\mathcal{Y}$, Remark 1 confirms that for all $\alpha\geq 0$, $\bm{\nu}+\alpha\bm{\eta}\in{\rm I\!R}^{n}\setminus\texttt{int}~{}\mathcal{V}$. ∎
@@
Figure 1: An illustration of Proposition 1. LABEL:sub@subfig:Conjugate-of-f1 Conjugate function $f^{*}_{1}$ of $f_{1}+\delta_{\mathcal{Y}_{1}}$. LABEL:sub@subfig:Conjugate-of-f2 Conjugate function $f^{*}_{2}$ of $f_{2}+\delta_{\mathcal{Y}_{2}}$. LABEL:sub@subfig:LevelSets-of-f1 Level sets of $f^{*}_{1}$ and associated set $\mathcal{V}$: $\nabla f_{1}^{*}$ over each ray remains intact, where $\mathcal{R}(\mathbf{v}_{1})$, $\mathcal{R}(\mathbf{v}_{2})$, and $\mathcal{R}(\mathbf{v}_{3})$ originates at $\mathbf{v}_{1}=[3\ 0.5]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{1}=[0\ 1]^{\mbox{\scriptsize T}}$, $\mathbf{v}_{2}=[0.3\ -4.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{2}=[0\ -1]^{\mbox{\scriptsize T}}$, and $\mathbf{v}_{3}=[4.5\ -4.5]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{3}=[1\ -1]^{\mbox{\scriptsize T}}$, respectively. LABEL:sub@subfig:LevelSets-of-f2 Level sets of $f^{*}_{2}$ and associated set $\mathcal{V}$: $\nabla f_{2}^{*}$ over each ray remains intact, where $\mathcal{R}(\mathbf{v}_{1})$, $\mathcal{R}(\mathbf{v}_{2})$, and $\mathcal{R}(\mathbf{v}_{3})$ originates at $\mathbf{v}_{1}=[-3\ 1.8]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{1}=[-5\ 3]^{\mbox{\scriptsize T}}$, $\mathbf{v}_{2}=[3.2\ 3.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{2}=[1\ 1]^{\mbox{\scriptsize T}}$, and $\mathbf{v}_{3}=[0.5\ -3.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{3}=[0\ -1]^{\mbox{\scriptsize T}}$, respectively.
@@
Consider the functions $f_{1}$ and $f_{2}$ in Example 1. The conjugate functions $f_{1}^{*}$ and $f_{2}^{*}$ of $f_{1}+\delta_{\mathcal{Y}_{1}}$ and $f_{2}+\delta_{\mathcal{Y}_{2}}$ are depicted in Figure 1LABEL:sub@subfig:Conjugate-of-f1 and Figure 1LABEL:sub@subfig:Conjugate-of-f2, respectively. The level sets of the conjugate functions are shown in Figure 1LABEL:sub@subfig:LevelSets-of-f1 and Figure 1LABEL:sub@subfig:LevelSets-of-f2 with their corresponding sets $\mathcal{V}$, cf.  Remark 1. The rays originated from some random points along which the gradients of the conjugate functions $f_{1}^{*}$ and $f_{2}^{*}$ remain constant are shown in the same plots.

### v3

@@ I-B Our Contribution
  • Application of FGOR to real-world problems: we explore how FGOR can be exploited when solving the global consensus problem [cf.  (33)], one of the most prevalent formulations in large-scale signal processing and machine learning applications. In particular, we show that by leveraging FGOR, the performance of the standard dual decomposition algorithm can be improved in terms of both speed of convergence and communication efficiency, cf.  § V, Algorithm 3, Lemma 2, and § VI.
  • Extensions to nonconvex and stochastic settings: we extend the established FGOR properties to nonconvex formulations and stochasic methods, cf.  Appendix D and Appendix E.
  • The numerical results highlight that FGOR can substantially improve the performance of existing stepsize methods [Polyak_Intro_Opt, goldstein_1962, armijo_1966, beck_2009, bello_2016, Asl_2020, khirirat_2023, Thinh_2018, Thinh_2021, zhor_1968_1, zhor_1968_2, goffin_1977, shor_2012, Damek_2018, zhu_2019, aybat_2019, chen_2019, wang_2021, krizhevsky_2012, He_2016, ge_2019, polyak_1969, hazan_2022, loizou_2021, wang_2023, barzilai_1988, dai_2002, burdakov_2019, Yura_2020] and, on average, outperform state-of-the-art splitting methods [condat_2013_splitting_methods, Boyd-Parikh-Chu-Peleato-Eckstein-2010, He_2021, Bai_2024] with respect to both convergence behavior and communication efficiency, cf.  § VI.
  • Lipschitzian properties of the dual function $g$: under less restricted assumptions, we establish that the gradients $\nabla g$ of $g$ are Lipschitz continuous, a more general Lipschitzian property than the existing results, cf.  Appendix F, Proposition 3, Corollary 3.
@@ I-C Notation
Normal font lowercase letters $x$, bold font lowercase letters $\mathbf{x}$, bold font uppercase letters $\mathbf{X}$, and calligraphic font $\mathcal{X}$ represent scalars, vectors, matrices, and sets, respectively. The set of real numbers, set of extended real numbers, set of real $n$-vectors, set of real $m\times n$ matrices, set of positive integers, and set of nonnegative integers are denoted by $\rm I\!R$}, $\overline{{\rm I\!R}}$, ${\rm I\!R}^{n}$, ${\rm I\!R}^{m\times n}$, $\mathbb{Z}_{+}$, and $\mathbb{Z}^{0}_{+}$, respectively. The boundary, the interior, the convex hull, and the closure of a set $\mathcal{X}$ are denoted by $\texttt{bnd}~\mathcal{X}$, $\texttt{int}~\mathcal{X}$, $\texttt{con}~\mathcal{X}$, and $\texttt{cl}~\mathcal{X}$, respectively. The domain of a function $f:{\rm I\!R}^{n}\to{\rm I\!R}$ is a subset of ${\rm I\!R}^{n}$ and is denoted by $\texttt{dom}~f$. The set of all the subgradients of a function $f$ at a point $\mathbf{x}\in{\rm I\!R}^{n}$ is denoted by $\partial f(\mathbf{x})$. The convex hull of a function $f:{\rm I\!R}^{n}\to{\rm I\!R}$ is denoted by $\texttt{con}~f$. The range and rank of a matrix $\mathbf{X}$ are denoted by $R(\mathbf{X})$ and $\texttt{rank}~(\mathbf{X})$, respectively. For a given matrix $\mathbf{X}$, $\mathbf{X}^{\mbox{\scriptsize T}}$ denotes the matrix transpose. The positive definite cone is denoted by $\mathbb{S}^{n}_{++}$. The $n\times n$ identity matrix and the $n$-vectors with all entries equal to one are denoted by $\mathbf{I}_{n}$ and $\mathbf{1}_{n}$, respectively.
@@ I-D Organization of the Paper
The rest of the paper is organized as follows. In § II, we investigate the FGOR characteristic of the conjugate function. The FGOR characteristic of the dual function is explored in § III. In § IV, the application of FGOR on the dual subgradient algorithm is discussed. Employing FGOR to solve the global consensus problem is discussed in § V. Numerical experiments are presented in § VI. Finally, we conclude the paper in § VII, followed by the appendices.
II FGOR Characteristic of the Conjugate
In this section, we demonstrate an appealing property of the conjugate function $f^{*}$ of $f$ [cf. problem (2)], that is, FGOR, which we will employ to devise our stepsize rules. Roughly speaking, we show that there is a specific region in the domain of $f^{*}$ such that for all points $\nu$ in the region, there is a ray originating from $\nu$ along which the gradients of $f^{*}$ remain constant. Besides, note that the dual function, $g$ is simply a restriction of $f^{*}$. As a result, further, we show that $g$ directly inherits the FGOR characteristic from $f^{*}$ if the restriction conforms to certain conditions. Then we leverage the FGOR characteristic of $g$ to design a precursory yet simple and communication-efficient stepsize rule that can be prepended with other existing ones for improved convergences.
@@ II FGOR Characteristic of the Conjugate
The function $f_{0}$ is lower semicontinuous, proper, and strictly convex. The set $\mathcal{Y}$ is convex and compact.
Note that since $f_{0}$ is strictly convex, $f^{*}$ is differentiable on ${\rm I\!R}^{n}$, and so is $g$, cf.  Appendix F, Lemma 4, Cor. 3. We further note that, the compact constraint set is not typically a restriction in practice. For example, it is a common approach in many machine learning problems that impose regularization as a constraint [singh_2018, gouk_2021]. It can be shown that the problem of convex regularized minimization is equivalent to its corresponding regularization-constrained formulation [bach_2012], cf.  Appendix B. Moreover, decision variables pertaining to feasible resources associated to engineering problems are usually finite [Boyd-Vandenberghe-04, Calafiore_2014]. Generalization to nonconvex $f_{0}$ is deferred to Appendix D.
@@
By the assumption above, one can exclude cases of infinite gradients at the boundary of $\mathcal{Y}$. The function $l=l_{0}+\delta_{\mathcal{Y}}$ in Example 7 [cf.  Appendix F] is a case where the assumption above breaks since the gradient of the function $l_{0}$ tends to $\infty$ or -$\infty$ as $\bar{y}$ tends to $1$ or 1-1, respectively.
Next, we present an important lemma and a few remarks that are useful in deriving the FGOR characteristic of $f^{*}$.
Lemma 1
@@
$\mathcal{V}=\texttt{cl}~\{\bm{{\nu}}\in\partial f_{0}(\mathbf{y})\ |\ \mathbf{y}\in\texttt{int}~\mathcal{Y}\}.$ (3)
Then $\bm{\nu}_{0}\in\texttt{int}~\mathcal{V}\iff\exists~\mathbf{y}\in\texttt{int}~\mathcal{Y}~\mbox{s.t.,}~\bm{\nu}_{0}\in\partial f(\mathbf{y})$.
Proof:
Let $\bm{\nu}_{0}\in\texttt{int}~\mathcal{V}$. Then, by the definition of $\mathcal{V}$, $\exists~\mathbf{y}\in\texttt{int}~\mathcal{Y}$ s.t., $\bm{\nu}_{0}\in\partial f_{0}(\mathbf{y})$. Since $f(\mathbf{y})=f_{0}(\mathbf{y})\,\forall~\mathbf{y}\in\texttt{int}~\mathcal{Y}$, $\bm{\nu}_{0}\in\partial f(\mathbf{y})$. Conversely, suppose that $\exists~\mathbf{y}\in\texttt{int}~\mathcal{Y}$ s.t., $\bm{\nu}_{0}\in\partial f(\mathbf{y})$. Again, since $f(\mathbf{y})=f_{0}(\mathbf{y})\,\forall~\mathbf{y}\in\texttt{int}~\mathcal{Y}$, $\bm{\nu}_{0}\in\partial f_{0}(\mathbf{y})$. Then, by the definition of $\mathcal{V}$, together with the strict convexity of $f_{0}$, $\bm{\nu}_{0}\in~\texttt{int}~\mathcal{V}$. ∎
We illustrate the set $\mathcal{V}$ [cf.  (3)] using the following example for clarity.
Example 1 (An Illustration of the Set $\mathcal{V}$)
@@
Then $\nabla f_{1}(\mathbf{y})=4\mathbf{y}$, where $\mathbf{y}\in\mathcal{Y}_{1}$. Thus, it is straightforward to see that the set $\mathcal{V}$ associated to $f_{1}$ is given by $\mathcal{V}=\{{\bm{\nu}\in{\rm I\!R}^{2}\ |\ -[0\ 4]^{\mbox{\scriptsize T}}\leq\bm{\nu}\leq[4\ 0]^{\mbox{\scriptsize T}}\}}$, cf.  Figure 1LABEL:sub@subfig:LevelSets-of-f1. Next, to illustrate the set $\mathcal{V}$ associated to $f_{2}$, we let $f_{21}=0.5\|\mathbf{y}-\mathbf{1}_{2}\|_{2}^{2}$ and $f_{22}=0.5\|\mathbf{y}+\mathbf{1}_{2}\|_{2}^{2}$. Then $\partial f_{2}(\mathbf{y})=\texttt{con}\{\nabla f_{2i}(\mathbf{y})\ |\ f_{2i}(\mathbf{y})=f_{2}(\mathbf{y}),i=1,2\}$, where $\mathbf{y}\in\mathcal{Y}_{2}$ since $f_{2}$ is the pointwise maximum of $f_{21}$ and $f_{22}$ [Boyd-EE364b-subgrad-22]. Therefore, it is easily seen that
$\partial f_{2}(\mathbf{y})=\begin{cases}\displaystyle\nabla f_{22}(y_{1},y_{2})\ ;\ \text{$y_{2}>-y_{1}$}\\ \displaystyle\nabla f_{21}(y_{1},y_{2})\ ;\ \text{$y_{2}<-y_{1}$}\\ \displaystyle\texttt{con}\{\nabla f_{21}(y_{1},y_{2}),\nabla f_{22}(y_{1},y_{2})\}\ ;\ \text{$y_{2}=-y_{1}$},\end{cases}$ (4)
[…]
@@
Suppose Assumption 1 and Assumption 2 hold. Then, for all $\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int~}\mathcal{V}$, there exists $\bm{\eta}\in{\rm I\!R}^{n}$ such that for all $\alpha\geq 0$, $\bm{\nu}+\alpha\bm{\eta}\in{\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}$ and $\nabla f^{*}(\bm{\nu}+\alpha\bm{\eta})$ is a constant vector that corresponds to some $\bar{\mathbf{y}}\in\texttt{bnd}~\mathcal{Y}$.
@@
Let $\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}$. Note that the equivalence established in Lemma 1 is equivalent to the contrapositive
$\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}\iff\forall~\mathbf{y}\in\texttt{int}~\mathcal{Y},\,\bm{\nu}\not\in\partial f(\mathbf{y}).$ (5)
However, since $\bm{\nu}\in\texttt{dom}~f^{*}$, it follows by the inversion rule for subgradient relations [Rockafellar-98, Prop. 11.3] that $\bm{\nu}\in\partial f(\mathbf{y})$ for some $\mathbf{y}\in\mathcal{Y}$. Thus, (5) yields
$\bm{\nu}\in{\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}\iff\exists~\mathbf{\bar{y}}\in\texttt{bnd}~\mathcal{Y},\,\mbox{s.t.,}~\bm{\nu}\in\partial f(\mathbf{\bar{y}}).$ (6)
Then, from (6) we have $\bm{\nu}\in\partial f(\mathbf{\bar{y}})$ for some $\mathbf{\bar{y}}\in\texttt{bnd}~\mathcal{Y}$, i.e.,
@@
where $\bar{\bm{\nu}}\in\partial f_{0}(\mathbf{\bar{y}})$ and $\bm{\eta}\in N_{\mathcal{Y}}(\mathbf{\bar{y}})$, cf.  Remark 2. Now consider the ray
@@
where the equality follows from (7). Since $\bar{\bm{\nu}}\in\partial f_{0}(\bar{\mathbf{y}})$ and $\bm{\eta}\in N_{\mathcal{Y}}(\bar{\mathbf{y}})$, $\bar{\bm{\nu}}+(1+\alpha)\bm{\eta}\in\partial f(\mathbf{\bar{y}})$ for all $\alpha\geq 0$, cf.  Remark 2. Therefore, [Rockafellar-98, Prop. 11.3], together with  (8), we have
@@
for all $\alpha\geq 0$. Moreover, since $\mathbf{\bar{y}}\in\texttt{bnd}~\mathcal{Y}$, equivalence (6) confirms that for all $\alpha\geq 0$, $\bm{\nu}+\alpha\bm{\eta}\in{\rm I\!R}^{n}\setminus\texttt{int}~\mathcal{V}$. ∎
@@
Figure 1: An illustration of Proposition 1. LABEL:sub@subfig:Conjugate-of-f1 $f^{*}_{1}$ of $f_{1}+\delta_{\mathcal{Y}_{1}}$. LABEL:sub@subfig:Conjugate-of-f2 $f^{*}_{2}$ of $f_{2}+\delta_{\mathcal{Y}_{2}}$. LABEL:sub@subfig:LevelSets-of-f1 Level sets of $f^{*}_{1}$ and associated set $\mathcal{V}$: $\nabla f_{1}^{*}$ over each ray remains intact, where $\mathcal{R}(\mathbf{v}_{1})$, $\mathcal{R}(\mathbf{v}_{2})$, and $\mathcal{R}(\mathbf{v}_{3})$ originates at $\mathbf{v}_{1}=[3\ 0.5]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{1}=[0\ 1]^{\mbox{\scriptsize T}}$, $\mathbf{v}_{2}=[0.3\ -4.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{2}=[0\ -1]^{\mbox{\scriptsize T}}$, and $\mathbf{v}_{3}=[4.5\ -4.5]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{3}=[1\ -1]^{\mbox{\scriptsize T}}$, respectively. LABEL:sub@subfig:LevelSets-of-f2 Level sets of $f^{*}_{2}$ and associated set $\mathcal{V}$: $\nabla f_{2}^{*}$ over each ray remains intact, where $\mathcal{R}(\mathbf{v}_{1})$, $\mathcal{R}(\mathbf{v}_{2})$, and $\mathcal{R}(\mathbf{v}_{3})$ originates at $\mathbf{v}_{1}=[-3\ 1.8]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{1}=[-5\ 3]^{\mbox{\scriptsize T}}$, $\mathbf{v}_{2}=[3.2\ 3.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{2}=[1\ 1]^{\mbox{\scriptsize T}}$, and $\mathbf{v}_{3}=[0.5\ -3.2]^{\mbox{\scriptsize T}}$ extending along $\bm{\eta}_{3}=[0\ -1]^{\mbox{\scriptsize T}}$, respectively.
@@
Consider $f_{1}$ and $f_{2}$ in Example 1. The conjugate functions $f_{1}^{*}$ and $f_{2}^{*}$ of $f_{1}+\delta_{\mathcal{Y}_{1}}$ and $f_{2}+\delta_{\mathcal{Y}_{2}}$ are depicted in Figure 1LABEL:sub@subfig:Conjugate-of-f1 and Figure 1LABEL:sub@subfig:Conjugate-of-f2, respectively. The level sets of the conjugate functions are shown in Figure 1LABEL:sub@subfig:LevelSets-of-f1 and Figure 1LABEL:sub@subfig:LevelSets-of-f2 with their corresponding sets $\mathcal{V}$, cf.  Lemma 1. The rays originated from some random points along which the gradients of $f_{1}^{*}$ and $f_{2}^{*}$ remain constant are shown in the same plots.